# Thermodynamic and energetic constraints on transition probabilities of small-scale quantum systems

### v2

@@ IV Conclusions
We have studied the transition probabilities of two-point measurement schemes for different initial thermal state occupations, in an otherwise generic quantum system. For such setups, we have proved two novel bounds that the transition probabilities (2) satisfy: i) The thermodynamic constraint (23a), stating that the response of the transition probabilities for different initial temperatures bounds from below the magnitude of the probabilities themselves multiplied by the heat dissipated while cooling or heating the system to these temperatures. ii) The energetic constraint (23b), which states that the response of the transition probabilities to temperature variations also bounds from below the transition probabilities weighted by the continuous change in the internal energy required to heat the system.
@@ Appendix A Work extraction and nonequilibrium free energy
$D[\hat{\rho}_{1}||\hat{\rho}_{2}]\equiv\text{Tr}\left\{\hat{\rho}_{1}(\log\hat{\rho}_{1}-\log{\hat{\rho}_{2}})\right\},$ (37)
@@
$\Delta S_{\text{S}}(\tau)=\beta\Delta U_{\text{S}}(\tau)+(\beta_{\text{B}}-\beta)U_{\text{S}}(\tau)+\log\left(\frac{Z_{\text{S}}(\tau)}{Z_{\text{S}}(0)}\right),$ (44)
with $Z_{\text{S}}(0)=\text{Tr}_{\mathrm{S}}\left\{\exp(-\beta\hat{H}_{\text{S}}(0))\right\}$ and $Z_{\text{S}}(\tau)=\text{Tr}_{\mathrm{S}}\left\{\exp(-\beta_{\text{B}}\hat{H}_{\text{S}}(\tau))\right\}$ being the partition functions at the inverse temperatures $\beta=T^{-1}$ and $\beta_{\text{B}}=T_{\text{B}}^{-1}$, respectively. We now recall that the internal energy of a thermal state is related to the partition function as $U_{\text{S}}(\tau)=-\frac{\partial}{\partial\beta_{\text{B}}}\log Z_{\text{S}}(\tau)$, and by using the concavity of $-\log Z_{\text{S}}$ as a function of the inverse temperature, Eq. (44) leads to the inequality
@@ Appendix B Sufficient condition for trivial constraints
Both thermodynamic and energetic costs, Eqs. (14) and (18) respectively, are positive by definition, and are lower-bounded by the response term $\mathcal{W}^{\text{(Resp)}}_{a\to b}$ in Eq. (13). Therefore, the thermodynamic and energetic constraints are nontrivial whenever $\mathcal{W}^{\text{(Resp)}}_{a\to b}>0$. In this Appendix we provide a sufficient condition for $\mathcal{W}^{\text{(Resp)}}_{a\to b}<0$, making the results of Eqs. (23a, 23b) trivial, and a sufficient condition for $\mathcal{W}^{\text{(Resp)}}_{a\to b}>0$, making the results of Eqs. (23a, 23b) nontrivial. Starting from the definition of the response contribution (13) and recalling that derivatives with respect to $\beta$ only act on the probabilities $p_{a}$, we consider
@@
$\partial_{\beta}p_{a}(\beta_{\text{h}})-\partial_{\beta}p_{a}(\beta_{\text{c}})\geq 0\quad\Rightarrow\quad\mathcal{W}^{\text{(Resp)}}_{a\to b}\leq 0,$ (53)
@@
$\partial_{\beta}p_{a}(\beta_{\text{h}})-\partial_{\beta}p_{a}(\beta_{\text{c}})\leq 0\quad\Rightarrow\quad\mathcal{W}^{\text{(Resp)}}_{a\to b}\geq 0.$ (54)

### v3

@@ IV Conclusions
We have studied the transition probabilities of two-point measurement schemes for different initial thermal state occupations, in an otherwise generic quantum system. For such setups, we have proved two novel bounds that the transition probabilities (2) satisfy: i) The thermodynamic constraint (12), stating that the response of the transition probabilities for different initial temperatures bounds from below the magnitude of the probabilities themselves multiplied by the heat dissipated while cooling or heating the system to these temperatures. ii) The energetic constraint (17), which states that the response of the transition probabilities to temperature variations also bounds from below the transition probabilities weighted by the continuous change in the internal energy required to heat the system.
@@ Appendix A Work extraction and nonequilibrium free energy
$D[\hat{\rho}_{1}||\hat{\rho}_{2}]\equiv\text{Tr}\left\{\hat{\rho}_{1}(\ln\hat{\rho}_{1}-\ln{\hat{\rho}_{2}})\right\},$ (37)
@@
$\Delta S_{\text{S}}(\tau)=\beta\Delta U_{\text{S}}(\tau)+(\beta_{\text{B}}-\beta)U_{\text{S}}(\tau)+\ln\left(\frac{Z_{\text{S}}(\tau)}{Z_{\text{S}}(0)}\right),$ (44)
with $Z_{\text{S}}(0)=\text{Tr}_{\mathrm{S}}\left\{\exp(-\beta\hat{H}_{\text{S}}(0))\right\}$ and $Z_{\text{S}}(\tau)=\text{Tr}_{\mathrm{S}}\left\{\exp(-\beta_{\text{B}}\hat{H}_{\text{S}}(\tau))\right\}$ being the partition functions at the inverse temperatures $\beta=T^{-1}$ and $\beta_{\text{B}}=T_{\text{B}}^{-1}$, respectively. We now recall that the internal energy of a thermal state is related to the partition function as $U_{\text{S}}(\tau)=-\frac{\partial}{\partial\beta_{\text{B}}}\ln Z_{\text{S}}(\tau)$, and by using the concavity of $-\ln Z_{\text{S}}$ as a function of the inverse temperature, Eq. (44) leads to the inequality
@@ Appendix B Sufficient condition for trivial constraints
Both thermodynamic and energetic costs, Eqs. (14) and (18) respectively, are positive by definition, and are lower-bounded by the response term $\mathcal{W}^{\text{(Resp)}}_{a\to b}$ in Eq. (13). Therefore, the thermodynamic and energetic constraints are nontrivial whenever $\mathcal{W}^{\text{(Resp)}}_{a\to b}>0$. In this Appendix we provide a sufficient condition for $\mathcal{W}^{\text{(Resp)}}_{a\to b}<0$, making the results of Eqs. (12, 17) trivial, and a sufficient condition for $\mathcal{W}^{\text{(Resp)}}_{a\to b}>0$, making the results of Eqs. (12, 17) nontrivial. Starting from the definition of the response contribution (13) and recalling that derivatives with respect to $\beta$ only act on the probabilities $p_{a}$, we consider
@@
$\partial_{\beta}p_{a}(\beta_{\text{h}})-\partial_{\beta}p_{a}(\beta_{\text{c}})\leq 0\quad\Rightarrow\quad\mathcal{W}^{\text{(Resp)}}_{a\to b}\leq 0,$ (53)
@@
$\partial_{\beta}p_{a}(\beta_{\text{h}})-\partial_{\beta}p_{a}(\beta_{\text{c}})\geq\,0\quad\Rightarrow\quad\mathcal{W}^{\text{(Resp)}}_{a\to b}\geq 0.$ (54)